OpenStudy (anonymous):

How do I apply the double angle formula to cos(450)?

OpenStudy (anonymous):

\[\cos 450 = \cos (360 + 90)= \cos 90\] Why do you need to use double angle formula?

OpenStudy (anonymous):

The question is asking me to find the exact value using the ouble angle formula. I did reduce to cos90 and did cos180=2cos^2(90)-1. But apparently i was wrong

OpenStudy (anonymous):

it was cos(90)

OpenStudy (anonymous):

So even applying double angle formula you get the same result?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so I'll skip that process next time thank you.

OpenStudy (anonymous):

welcome :)